OpenStudy (anonymous):

Please help! I will give medals and fan for help! :) Find the constant of variation for the relationship f(x)= 50x.

OpenStudy (anonymous):

50?

OpenStudy (anonymous):

Idk can someone explain this to me? please?

OpenStudy (mathstudent55):

If you take x = 1, f(1) = 50(1) = 50. For x = 2, f(2) = 50(2) = 100. Notice that each time, x was multiplied by 50 to become f(x). 50/1 = 50 100/2 = 50

OpenStudy (anonymous):

so the constant is f(x)?

OpenStudy (mathstudent55):

The constant of variation is 50 because that is the number you always get when you divide the y-coordinate by the x-coordinate of any point in this function.

OpenStudy (anonymous):

The Constant Variation would be 50.

OpenStudy (anonymous):

oh ok thanks mind if I tag you in other questions for more help?